Find the value of r so that the line through (4,5) and (r,3) has a slope of 2/3
Answer by mananth(16949)

The slope of line passing through (x1,y1),(x2,y2) = (y2-y1)/(x2-x1)

(3-5)/(r-4)=2/3
2(r-4)=-6
(r-4)=-3
r=1